# Multilinguality Does not Make Sense: Investigating Factors Behind Zero-Shot Transfer in Sense-Aware Tasks

### v1

@@ Abstract
Cross-lingual transfer allows models to perform tasks in languages unseen during training and is often assumed to benefit from increased multilinguality. In this work, we challenge this assumption in the context of two underexplored, sense-aware tasks: polysemy disambiguation and lexical semantic change. Through a large-scale analysis across 28 languages, we show that multilingual training is neither necessary nor inherently beneficial for effective transfer. Instead, we find that confounding factors—such as fine-tuning data composition and evaluation artifacts—better account for the perceived advantages of multilinguality. Our findings call for more rigorous evaluations in multilingual NLP. We release fine-tuned models and benchmarks to support further research, with implications extending to low-resource and typologically diverse languages.
Multilinguality Does not Make Sense: Investigating Factors Behind Zero-Shot Transfer in Sense-Aware Tasks
Roksana Goworek1 Haim Dubossarsky1,2 1 Queen Mary University of London 2 Language Technology Lab, University of Cambridge {r.goworek, h.dubossarsky}@qmul.ac.uk
@@ 1 Introduction
Cross-lingual transfer enables models to leverage knowledge acquired in one language to perform tasks in another, and requires pre-training on multilingual corpora (e.g., Wu and Dredze, 2019; Ponti et al., 2018). This ability underpins many of today’s advances in multilingual NLP and has been evaluated across a wide range of tasks - from syntactic parsing and POS tagging to more complex, semantics-driven tasks like question answering, language inference, and paraphrasing (see Philippy et al. (2023) for a comprehensive review).
Polysemy disambiguation - despite being foundational to linguistic meaning and long-standing challenges in NLP (Navigli, 2009; Bevilacqua et al., 2021) - received little attention in transfer learning research. Polysemy poses unique challenges to transfer due to its inherently language-specific nature (Rzymski et al., 2020). For example, while the English word mole denotes ‘a small burrowing mammal’ and ‘a skin blemish,’ its Hindi counterpart
(til) refers to the latter sense but also means ‘sesame seed’. Such non-aligned senses across languages complicate direct transfer, which makes it an ideal testbed for evaluating the true limits of cross-lingual generalization.
A common assumption is that multilinguality itself, set by a model’s exposure to multiple languages, is key for cross-lingual transfer. While it is evident that models must support both source and target languages, the extent to which broader multilinguality facilitates transfer remains unclear and understudied. Does access to more languages intrinsically improve transfer, or do other factors Philippy et al. (2023), like language similarity, or training confounds, drive the observed gains?
@@
While our findings may seem narrowly focused, they shed light on broader issues central to NLP, with implications for low-resource and domain-specific settings for which relying on broader multilingual models is not an option, as well as the wider research community. We release the best-performing models to support further analysis.
@@ 2 Related Work
Cross-lingual transfer studies have expanded from tasks like syntactic parsing, POS tagging, and semantic classification Wu and Dredze (2019); Choenni et al. (2023a); Pires et al. (2019); de Vries et al. (2022) to more complex ones such as NER, NLI Dolicki and Spanakis (2021); Srinivasan et al. (2021), and more recently QA, paraphrasing, and sentiment analysis Lauscher et al. (2020); Ahuja et al. (2023); Choenni et al. (2023b); Wang et al. (2023). Most studies rely on pretrained multilingual models like XLM-R and mBERT, typically fine-tuned on a source language and evaluated on target languages.
@@
Despite extensive work, the role of multilinguality itself is rarely tested directly and is often assumed to be beneficial. A notable exception is Shaham et al. (2024), who compare PaLM 2 fine-tuned on monolingual vs. multilingual data for instruction tuning, and find modest multilingual exposure aids transfer, while too much can degrade it. However, their analysis is limited to a single model and task, leaving open questions about pretraining and generalizability. Choenni et al. (2023b) show source language examples can influence target predictions, but without explicitly controlling for multilinguality, offering only indirect evidence. Chang et al. (2024) directly manipulated multilinguality by pretraining over 10,000 models across 250 languages, but evaluated only with perplexity, omitting standard tasks and making comparisons difficult and replication impractical.
@@
Exceptions for polysemy are few, Raganato et al. (2020) reported transfer to 12 languages, but use only English as a source language, while Dairkee and Dubossarsky (2024) found near-zero transfer to Hindi, questioning the feasibility of zero-shot transfer in low-resource settings. Arefyev et al. (2021) showed that training on polysemy disambiguation generalizes to semantic change detection, linking the two tasks. Cassotti et al. (2023) fine-tuned models on multilingual polysemy datasets, achieving state-of-the-art results Schlechtweg et al. (2020) and strong transfer to unseen languages Periti and Tahmasebi (2024), though none of these works isolate the effect of multilinguality itself.
Overall, while multilingual transfer has been widely studied, its direct contribution - especially in lexically focused tasks like polysemy and LSC - remains underexplored. This work addresses that gap by systematically testing the role of multilinguality in sense-aware transfer. By manipulating multilingual conditions, we clarify when and how multilinguality supports cross-lingual generalization, helping to resolve prior conflicting findings and inform future research on transfer learning.
@@ 3 Methods
In this study we set to isolate multilinguality from confounds, enabling a clear assessment of its independent impact on zero-shot transfer in polysemy disambiguation and LSC tasks. We conduct a large-scale evaluation of zero-shot and full-shot cross-lingual transfer performance across 28 languages, focusing on the unique contribution of multilingual training. To achieve this, we implement an experimental framework that systematically controls for potential confounds, ensuring that observed effects are attributed to multilinguality rather than artifacts of training data size or pretraining exposure.
@@ 3.2 Data
MCL, XL and Hindi datasets, which together span 18 languages, were used for training and evaluation, while AM2iCO and LSCD, spanning 14 and 7 languages, respectively, were used only for evaluation.111AM2iCO uses English as a pivot language therefore is biased toward it; LSCD is not in the WiC format. As some languages have only development and test data, we followed Cassotti et al. (2023) who sampled 75% of the development data of each language to enable training on these languages (keeping the remaining 25% for setting hyperparameters), using his exact train-dev-test splits. All WiC datasets are class-balanced, setting the chance baseline at 50%. German, French and English are over represented in the dataset relative to other languages (see Figure 4).
@@
XL by Raganato et al. (2020) used WordNet of Bulgarian, Chinese, Croatian, Danish, Dutch, English, Estonian, Japanese, Korean and Farsi, filtering out fine-grained senses. French, German and Italian used Wiktionary. Mean human accuracy was 80%, and varied between 74% for German, 87% for Danish and 97% for Farsi.
@@
AM2iCO by Liu et al. (2021) has English paired with 14 target languages. Compiled from Wikipedia Dumps of each language by selecting words with at least two different Wikipedia pages that shows ambiguity in both the target language and English. Overall human accuracy was 90.6%, with an inter-annotator agreement of 88.4%.
Lexical Semantic Change Detection (LSCD) covers seven languages from different sources: English, German, Latin, and Swedish from Schlechtweg et al. (2020), Spanish from Zamora-Reina et al. (2022), Chinese from Chen et al. (2023), and Norwegian from Kutuzov et al. (2022).222Norwegian1, Norwegian2, are two corpora pairs, comparing the words from four time periods. For each language and target word, equal number of sentences were sampled from corpus 1 (historical) and corpus 2 (modern).
@@ 3.3 Multilingual Base Models
XLM-R Conneau et al. (2020) and mBERT Devlin et al. (2019) are pretrained on 100 and 104 languages, respectively, covering all languages in our study, which enables us to train and evaluate on all languages in zero- and full-shot transfer. Both models are commonly used in multilingual research, XLM-R was extensively used in the context of WiC and LSC (Raganato et al., 2020; Cassotti et al., 2023; Dairkee and Dubossarsky, 2024), providing a strong baseline for comparison.
@@
LLaMA3-8B-Instruct Grattafiori et al. (2024); AI@Meta (2024) is a state-of-the-art instruction-tuned language model with 8 billion parameters, significantly larger than the other models. Pretraining data is not public, and was assumed to include major NLP datasets Sainz et al. (2023), like WiC.
MuRIL (Multilingual Representations for Indian Languages) by Khanuja et al. (2021) is pretrained on 16 Indian languages + English, which, along with Hindi, are the only overlapping languages with our study. It was used to test an edge-case of zero-shot transfer (see §4.2).
@@ 3.4 Model Training and Testing
To control for the different training sizes across languages, fixed-size versions of datasets were created for the multi, english, french and german datasets by randomly subsampling 8,750 examples from their training sets.333Other languages were too small to subsample.
Finetuning
All models were finetuned in the same way, selectively focusing on different subsets of the WiC datasets.Following Cassotti et al. (2023) and after a failed pilot study with cross-encoders, a Siamese bi-encoder was used to generate two distinct vector representations (embeddings) for the two usages of the target word in the two sentences.444 Training hyperparameters and model sizes are reported in Appendix 6. The model outputs the cosine distance between the output embeddings of the two inputs, and, in order to adapt it to the binary nature of WiC, a threshold is applied to decide if the words are classified as having the same sense. The model is trained to update its parameters (i.e., embeddings) to maximize this distance when the target appears in different meanings (label 0) and minimize it when the meanings are the same in the two sentences (label 1), by minimizing contrastive loss. During training, as well as inference, special tokens, <t> and </t>, are placed around the target word in each sentence to signal what word the model should focus on. See Appendix I.2 for LLaMA’s finetuning details.
@@ 4.1 Multilinguality Effects and Confounds
To rigorously assess the role of multilinguality in zero-shot transfer, we conduct three types of controlled comparisons, ensuring that fine-tuning size is held constant and distinguishing between full-shot and true zero-shot transfer scenarios. We provide a comprehensive evaluation of four multilingual base-models (§3.3) on 28 languages, first examining whether multilingual training offers an advantage and then systematically accounting for key confounding factors. We summarize the average accuracies of the models in Figure 1 and present the true zero-shot transfer comparison in Table 2. The full results are reported in Appendix G. Due to space constraints, except for Figure 1, we show and discuss XLM-R results in the main text, only briefly discussing other models while providing their full results in the Appendices.
@@
multi is fine-tuned on an order of magnitude more data than any other model (see Figure 4), which gives it an unfair advantage over all monolingual models. We therefore fix the size of the training datasets across all models (§3.4) to enable a fair comparison, and repeat the same analysis under these controlled conditions. The results in Figure 1, and especially the comparison between Full F.T and Fixed F.T (blue and orange bars, respectively) which are summarized in Table 1, show that the performance of the multi model drops much more than mono English when training size is controlled, perhaps due to a larger relative drop in training size, making mono English the best model across most datasets. Similar drops are observed for mBERT, BLOOM, and LLaMA (Tables 7-9), which for the latter largely diminished the advantage it had in Full F.T . Despite its drop in the Fixed F.T condition, LLaMA seems to benefit from multilingual fine-tuning more than other models. We attribute this to its nature as a generative model and pretraining uncertainty which may include data contamination unwanted for our experiments Sainz et al. (2023). Additionally, LLaMA relies on English prompts with explicit task instructions, which showed improved performance, particularly for hindi and chinese, where training data is more limited. Overall, this contrast (blue and orange) undermine the notion that training on a multilingual dataset inherently improves transfer. Instead, the advantage originally observed for multi in the Full F.T may largely be due to more training data (122.4k examples for multi cf. 54.7k for German, 46.1k for French, and 15.1k for English; see Table 4). This finding does not imply that training on more examples is not a good strategy to improve transfer, which it clearly is, only that attributing the performance gains to multilinguality is flawed.
@@
If multilinguality provides a novel sense-understanding, then not only should its performance improve, but also its errors should differ from monolingual models that lack exposure to such linguistic diversity. Instead, Figure 6 shows greater overlap of errors between multi and mono models, which even increases for Fixed F.T, supporting a notion that multi is simply another mono model (see formulation of alignment measures in Appendix E). This provides converging evidence that supports previous findings, further undermining the importance of multilinguality.
@@ 4.2 Underlying Factors of Successful Transfer
To test this hypothesis, and understand what drives variation in transfer results, we computed Pearson correlations between languages’ log-transformed pretraining sizes in models for which pretraining sizes were available (XLM-R, mBERT, BLOOM) and their corresponding accuracies on the four WiC datasets and on LSCD.555Pretraining sizes are provided in Appendix 13. For BLOOM, we removed languages not in its pretraining.
Figure 2 shows strong correlations between the pretraining size of a language and its accuracy as a target language. We attribute BLOOM’s poor performance to its unusual pretraining language distribution that focuses on low-resource languages. The lack of correlation in XL remains unclear and requires further investigation (see §5 for discussion).
@@
We tested the link between linguistic similarity and transferability using syntactic similarity scores Littell et al. (2017), which quantify similarity based on shared syntactic features, as is standard in NLP Philippy et al. (2023). Pearson correlations between similarity scores and model accuracies were computed across target language pairs. Syntactic similarity showed some correlation with zero-shot performance, the effect was weaker and less consistent than that of pretraining size (see Appendix §5). Moreover, we found that syntactic similarity is highly correlated with pretraining size, further undermining its contribution.
@@
While the correlation analysis provides useful insights, it remains slightly inconclusive. A key limitation is that we cannot pretrain models and directly manipulate this variable for a controlled experiment. We also lack sufficient fine-tuning data for languages that were less prominent during pretraining of our models, making it difficult to determine whether poor model performance stems from limited fine-tuning or limited exposure during pretraining.
To counter this, we analyzed transfer in an edge case scenario. We used MuRIL (§3.3), a model pretrained exclusively on 16 Indian languages and English. Evaluating zero-shot transfer performance of models that are fine-tuned on languages entirely absent from MuRIL’s pretraining corpus, can offer insight on the role of pretraining size in transfer.
Figure 3 shows graded transfer effects associated with the pretraining size, supporting the correlation analysis above. Hindi, being the dominant language in MuRIL (despite less data than English, it benefits from exposure to 16 Indic languages, including 2–4 with the same script and 9 with lexical overlap), enjoys the largest zero-shot transfer regardless of the fine-tuning language (tallest blue bars across all different models, not considering the full-shot condition in white), followed by English (yellow bars). Second, we observe substantial transfer from French, German and even Chinese on Hindi, languages MuRIL was not pretrained on (see §5 for further discussion). Third, MuRIL was able to learn and perform well on German, despite being absent during its pretraining stage.

### v2

# Multilinguality Does not Make Sense: Investigating Factors Behind Zero-Shot Cross-Lingual Transfer in Sense-Aware Tasks

@@ Abstract
Cross-lingual transfer allows models to perform tasks in languages unseen during training and is often assumed to benefit from increased multilinguality. In this work, we challenge this assumption in the context of two underexplored, sense-aware tasks: polysemy disambiguation and lexical semantic change. Through a large-scale analysis across 28 languages, we show that multilingual training is neither necessary nor inherently beneficial for effective transfer. Instead, we find that confounding factors, such as fine-tuning data composition and evaluation artifacts, can better account for the perceived advantages of multilinguality. Our findings call for more rigorous evaluations in multilingual NLP, and more nuanced and sensible choice of models for transfer. We release fine-tuned models and benchmarks to support further research, with implications extending to low-resource and typologically diverse languages.
Multilinguality Does not Make Sense: Investigating Factors Behind Zero-Shot Cross-Lingual Transfer in Sense-Aware Tasks
Roksana Goworek1 Haim Dubossarsky1,2,3 1 Queen Mary University of London 2 The Alan Turing Institute 3 University of Cambridge {r.goworek, h.dubossarsky}@qmul.ac.uk
@@ 1 Introduction
Cross-lingual transfer enables multilingual pretrained models to leverage knowledge acquired in one language to perform tasks in another (e.g., Wu and Dredze, 2019; Ponti et al., 2018). This ability underpins many of today’s advances in multilingual NLP and has been evaluated across a wide range of tasks - from syntactic parsing and POS tagging to more complex, semantics-driven tasks like question answering, language inference, and paraphrasing (see Philippy et al. (2023) for a review).
Polysemy disambiguation, despite being foundational to linguistic meaning and a long-standing challenge in NLP (Navigli, 2009; Bevilacqua et al., 2021), has received relatively little attention in transfer learning research. Polysemy poses unique challenges to transfer due to its inherently language-specific nature (Rzymski et al., 2020). For example, while the English word mole denotes ‘a small burrowing mammal’ and ‘a skin blemish,’ its Hindi counterpart
(til) refers to the latter sense but also means ‘sesame seed’. Such non-aligned senses across languages complicate direct transfer, making it an ideal testbed for evaluating the true limits of cross-lingual generalization.
A common assumption is that multilinguality itself, set by a model’s exposure to multiple languages, is key for cross-lingual transfer. While it is evident that models must support both source and target languages, the extent to which broader multilinguality facilitates transfer remains unclear and understudied. Does access to more languages intrinsically improve transfer, or do other factors, such as language similarity or training setup, drive the observed gains?
@@
While our findings may seem narrowly focused, they illuminate broader issues central to NLP, with implications for low-resource and domain-specific settings for which relying on broader multilingual models and corpora is not an option, as well as the wider research community. We release the best-performing models to support further analysis.111Best-performing mono and multi models are available in the collection Multilinguality Does not Make Sense.
@@ 2 Related Work
Cross-lingual transfer studies have expanded from tasks like syntactic parsing, POS tagging, and semantic classification Wu and Dredze (2019); Choenni et al. (2023a); Pires et al. (2019); de Vries et al. (2022) to more complex ones such as NER, NLI Dolicki and Spanakis (2021); Srinivasan et al. (2021), and more recently QA, paraphrasing, and sentiment analysis Lauscher et al. (2020); Ahuja et al. (2023); Choenni et al. (2023b); Wang et al. (2023). Most studies rely on pretrained multilingual models like XLM-R and mBERT.
@@
Despite extensive work, the role of multilinguality itself is rarely tested directly and is often assumed to be beneficial. A notable exception is Shaham et al. (2024), who compare PaLM 2 fine-tuned on monolingual versus multilingual data for instruction tuning, and find that modest multilingual exposure aids transfer, while too much can degrade it. However, their analysis is limited to a single model and task, leaving open questions about pretraining and generalizability. Choenni et al. (2023b) show that source language examples can influence target predictions, but without explicitly controlling for multilinguality, offering only indirect evidence. Chang et al. (2024) directly manipulated multilinguality by pretraining over 10,000 models across 250 languages, but only evaluated using perplexity with no standard tasks, making comparisons difficult and replication impractical.
@@
Exceptions for polysemy are few. Raganato et al. (2020) reported transfer to 12 languages, but used only English as a source language, while Dairkee and Dubossarsky (2024) found near-zero transfer to Hindi, raising questions about the feasibility of zero-shot transfer in low-resource settings. In contrast, Goworek et al. (2025) recently showed notable zero-shot transfer from English to 10 low-resource languages, highlighting the need for further investigation.
With regard to LSC, Arefyev et al. (2021) showed that training on polysemy disambiguation generalizes to semantic change detection, linking the two tasks. Cassotti et al. (2023) followed up and fine-tuned models on multilingual polysemy datasets, achieving state-of-the-art results Schlechtweg et al. (2020) and strong transfer to unseen languages Periti and Tahmasebi (2024).
None of these works studied multilinguality itself. A notable exception is Berend (2022), who explicitly examined the role of multilinguality in word sense disambiguation transfer. However, their study investigated the role of multilinguality only at the pretraining stage, rather than during fine-tuning. This is a less practical perspective given the ubiquitous use of multilingual models in NLP today, where many languages lack high quality monolingual models.
Overall, while multilingual transfer has been widely studied, the direct contribution of multilinguality in fine-tuning remains underexplored, especially in lexically focused tasks like polysemy and LSC. This work addresses that gap by systematically testing the role of multilinguality in sense-aware transfer. By manipulating multilingual conditions, we clarify when and how multilinguality supports cross-lingual generalization, helping to resolve prior conflicting findings and inform future research on transfer learning.
@@ 3 Methods
In this study, we set to isolate multilinguality from confounds, enabling a clear assessment of its independent impact on zero-shot transfer in polysemy disambiguation and LSC tasks. We conduct a large-scale evaluation of zero-shot and full-shot cross-lingual transfer performance across 28 languages, focusing on the unique contribution of multilingual training. To achieve this, we implement an experimental framework that systematically controls for potential confounds, ensuring that observed effects are attributed to multilinguality rather than artifacts of training data size or pretraining exposure.
@@ 3.2 Data
MCL, XL and Hindi datasets, which together span 18 languages, were used for training and evaluation, while AM2iCO and LSCD (LSC Detection), spanning 14 and 7 languages, respectively, were used only for evaluation.222AM2iCO uses English as a pivot language therefore is biased toward it; LSCD is not in the WiC format. As some languages have only development and test data, we followed Cassotti et al. (2023) who sampled 75% of the development data of each language to enable training on these languages (keeping the remaining 25% for setting hyperparameters), using their exact train-dev-test splits. All WiC datasets are class-balanced, setting the chance baseline at 50%. German, French and English are overrepresented in the dataset relative to other languages (see Figure 4).
@@
XL by Raganato et al. (2020) used WordNet of Bulgarian, Chinese, Croatian, Danish, Dutch, English, Estonian, Japanese, Korean and Farsi, filtering out fine-grained senses. French, German and Italian used Wiktionary. The reported mean human accuracy was 80%, and varied between 74% for German, 87% for Danish and 97% for Farsi.
@@
AM2iCO by Liu et al. (2021) has English paired with 14 target languages. Compiled from Wikipedia Dumps of each language by selecting words with at least two different Wikipedia pages that show ambiguity in both the target language and English. Overall human accuracy was 90.6%, with an inter-annotator agreement of 88.4%.
Lexical Semantic Change Detection (LSCD) covers seven languages from different sources: English, German, Latin, and Swedish from Schlechtweg et al. (2020), Spanish from Zamora-Reina et al. (2022), Chinese from Chen et al. (2023), and Norwegian from Kutuzov et al. (2022).333Norwegian1, Norwegian2, are two corpus pairs, comparing the words from four time periods. For each language and target word, an equal number of sentences were sampled from corpus 1 (historical) and corpus 2 (modern).
@@ 3.3 Multilingual Base Models
XLM-R-large Conneau et al. (2020) and mBERT Devlin et al. (2019) are pretrained on 100 and 104 languages, respectively, covering all languages in our study, which enables us to train and evaluate on all languages in zero- and full-shot transfer. Both models are commonly used in multilingual research. XLM-R was extensively used in the context of WiC and LSC (Raganato et al., 2020; Cassotti et al., 2023; Dairkee and Dubossarsky, 2024), providing a strong baseline for comparison.
@@
LLaMA3-8B-Instruct Grattafiori et al. (2024); AI@Meta (2024) is a decoder-only instruction-tuned language model with 8 billion parameters, significantly larger than the other models. Pretraining data is not public, and is assumed to include major NLP datasets Sainz et al. (2023), like WiC.
MuRIL (Multilingual Representations for Indian Languages) by Khanuja et al. (2021) is pretrained on 16 Indian languages + English, which, along with Hindi, are the only languages overlapping with our study. It was used to test an edge case of zero-shot transfer (see §4.2).
@@ 3.4 Model Training and Testing
To control for the different training sizes across languages, fixed-size versions of datasets were created for the multi, english, french and german datasets by randomly subsampling 8,750 examples from their training sets.444Other languages had too little data to subsample.
Finetuning Details
All encoder models were fine-tuned in the same way, selectively focusing on different subsets of the WiC datasets. Following Cassotti et al. (2023) and after a failed pilot study with cross-encoders, a Siamese bi-encoder was used to generate two distinct vector representations (embeddings) for the two usages of the target word in the two sentences. The model outputs the cosine distance between the output embeddings of the two inputs, and, in order to adapt it to the binary nature of WiC, a threshold is applied to decide if the words are classified as having the same sense. The model is trained to update its parameters (i.e., embeddings) to maximize this distance when the target appears in different meanings (label 0) and minimize it when the meanings are the same in the two sentences (label 1), by minimizing contrastive loss. During training, as well as inference, special tokens, <t> and </t>, are placed around the target word in each sentence to signal what word the model should focus on.
To preserve comparability and avoid cherry-picking, we fixed a single hyperparameter configuration and did not perform per-model/per-language tuning or multi-seed sweeps, which may understate best-case performance and limit our characterization of variance across runs. For more discussion of this decision, see Appendix E.555Training hyperparameters and model sizes are in Table 6.
To verify that performance is significantly improved by fine-tuning rather than being the result of inherent sense-distinction ability of contextualized embeddings, we conduct no-fine-tuning experiments with frozen pretrained models. The results corroborate this and can be found in Appendix I. For details of LLaMA’s fine-tuning (same data, different procedure), see Appendix L.1.
@@ 4.1 Multilinguality Effects and Confounds
To rigorously assess the role of multilinguality in zero-shot transfer, we conduct three types of controlled comparisons, ensuring that fine-tuning size is held constant and distinguishing between full-shot and true zero-shot transfer scenarios. We provide a comprehensive evaluation of four multilingual base-models (§3.3) on 28 languages, first examining whether multilingual training offers an advantage and then systematically accounting for key confounding factors. We summarize the average accuracies of the models in Figure 1 and present the true zero-shot transfer comparison in Table 2. Due to space constraints, except for Figure 1, we show and discuss XLM-R results in the main text, only briefly discussing other models while providing their full results in Appendix H.
@@
multi is fine-tuned on an order of magnitude more data than any other model (see Figure 4), which gives it an unfair advantage over all monolingual models. We therefore fix the size of the training datasets across all models (§3.4) to enable a fair comparison, and repeat the same analysis under these controlled conditions. The results in Figure 1, and especially the comparison between Full F.T and Fixed F.T (blue and orange bars, respectively) which are summarized in Table 1, show that the performance of the multi model drops much more than mono English when training size is controlled, perhaps due to a larger relative drop in training size, making mono English the best model across most datasets. Similar drops are observed for mBERT, BLOOM, and LLaMA (Tables 7-9), which for the latter largely diminished the advantage it had in Full F.T. Despite its drop in the Fixed F.T condition, LLaMA seems to benefit from multilingual fine-tuning more than other models. We attribute this to its nature as a generative model and pretraining uncertainty which may include data contamination unwanted for our experiments Sainz et al. (2023). Additionally, LLaMA relies on English prompts with explicit task instructions, which showed improved performance, particularly for hindi and chinese, where training data is more limited. Overall, this contrast (blue and orange) undermines the notion that training on a multilingual dataset inherently improves transfer. Instead, the advantage originally observed for multi in the Full F.T may largely be due to more training data (122.4k examples for multi cf. 54.7k for German, 46.1k for French, and 15.1k for English; see Table 4). This finding does not imply that training on more examples is not a good strategy to improve transfer, which it clearly is, only that attributing the performance gains to multilinguality is flawed.
@@
If multilinguality equips models with novel sense-understanding, then not only should its performance improve, but also its errors should differ from mono models that lack exposure to such linguistic diversity. Instead, Figure 6 shows greater overlap of errors between multi and mono models, which even increases for Fixed F.T, supporting a notion that multi is simply another mono model (see formulation of alignment measures in Appendix F). This provides converging evidence that supports previous findings, further undermining the importance of multilinguality.
@@
The z.s. multi⊗= row presents the results of five different multi⊗= models (each trained on all available languages excluding the target language, specified by the column, respectively). The {lang}= rows show the results of the specified monolingual model on the test set specified by the column. Grey cells mark full-shot conditions for monolingual models.
@@ 4.2 Underlying Factors of Successful Transfer
To test this hypothesis, and understand what drives variation in transfer results, we computed Pearson correlations between languages’ log-transformed pretraining sizes in models for which pretraining sizes were available (XLM-R, mBERT, BLOOM) and their corresponding accuracies on the four WiC datasets and on LSCD.666Pretraining sizes are provided in Appendix 13. For BLOOM, we removed languages not in its pretraining.
Figure 2 shows strong correlations between a language’s pretraining size and its accuracy as a target language. We attribute BLOOM’s poor correlations to its unusual pretraining language distribution that focuses on low-resource languages. The lack of correlation in XL remains unclear and requires further investigation (see §5 for discussion).
@@
We tested the link between linguistic similarity and transferability using syntactic similarity scores Littell et al. (2017), which quantify similarity based on shared syntactic features, as is standard in NLP Philippy et al. (2023). Pearson correlations between similarity scores and model accuracies were computed across target language pairs. Syntactic similarity showed some correlation with zero-shot performance, the effect was weaker and less consistent than that of pretraining size (see Appendix 5). Moreover, we found that syntactic similarity is highly correlated with pretraining size, further undermining its contribution.
@@
While the correlation analysis provides useful insights, it remains slightly inconclusive. A key limitation is that we cannot pretrain models to directly manipulate this variable for a controlled experiment. We also lack sufficient fine-tuning data for languages that were less prominent during pretraining, making it difficult to determine whether poor model performance stems from limited fine-tuning or limited exposure during pretraining.
To counter this, we analyzed transfer in an edge case scenario. We used MuRIL (§3.3), a model pretrained exclusively on 16 Indian languages and English. Evaluating zero-shot transfer performance of models that are fine-tuned on languages entirely absent from MuRIL’s pretraining corpus can offer insight on the role of pretraining size in transfer.
Figure 3 shows graded transfer effects associated with the pretraining size, supporting the correlation analysis above. Hindi, being the dominant language in MuRIL (despite less data than English, it benefits from exposure to 16 Indic languages, including 2–4 with the same script and 9 with lexical overlap), enjoys the largest zero-shot transfer regardless of the fine-tuning language (tallest blue bars across all different models, not considering the full-shot condition in white), followed by English (yellow bars). Second, we observe substantial transfer from French, German and even Chinese on Hindi, languages MuRIL was not pretrained on (see §5 for further discussion). Third, MuRIL was able to learn and perform well on German, despite it being absent from its pretraining stage.